### v2

# On the topological ranks of Banach -algebras associated with groups of subexponential growth

Abstract
[…]
1 Introduction
The (left) topological stable rank of a unital Banach algebra $\mathfrak{B}$ was introduced by Rieffel [Ri83] as the smallest $n\in\mathbb{Z}^{+}$ such that
${\rm Lg}_{n}(\mathfrak{B})=\left\{(b_{1},\ldots,b_{n})\in\mathfrak{B}^{n}\mid\mathfrak{B}=\mathfrak{B}b_{1}+\ldots+\mathfrak{B}b_{n}\right\}$
is dense in $\mathfrak{B}^{n}$. The idea was to produce a noncommutative notion of dimension similar to both the covering dimension for compact spaces and the Bass stable rank. This idea was then applied very successfully to the study of K-theory for $C^{*}$-algebras [Ri83b] and motivated the notion of real rank. Indeed, the real rank of a unital $C^{*}$-algebra $\mathfrak{B}$, as introduced by Brown and Pedersen [BP91], is the smallest $n\in\mathbb{N}$ such that
${\rm Lg}_{n+1}(\mathfrak{B})_{\rm sa}=\left\{(b_{0},\ldots,b_{n})\in\mathfrak{B}_{\rm sa}^{n+1}\mid\mathfrak{B}=\mathfrak{B}b_{0}+\ldots+\mathfrak{B}b_{n}\right\}$
is dense in $\mathfrak{B}_{\rm sa}^{n+1}$.
Studying these noncommutative notions of dimension became a very interesting problem for $C^{*}$-algebraists and motivated many works within the area. This is particularly true in the case of group $C^{*}$-algebras and crossed products. Just to mention some examples, the topological stable rank has been studied for group $C^{*}$-algebras in [SuTa95, DyHaRo97], whereas the real rank of group $C^{*}$-algebras has been studied in [Ka93, ArKa12, Mo23]. In the setting of crossed products, we can also find the works [Pu90b, Pu90, CP05].
There are also many works on topological ranks that stay outside the realm of $C^{*}$-algebras (see [Ba98, Ba99, DaJi09]) and this will be our case too. Indeed, we will be interested in the topological ranks of (generalized) convolution algebras and $L^{p}$-operator algebras, as introduced by Christopher Phillips [CP12, CP13].
The study of $L^{p}$-operator algebras has attracted much attention in the last decade (see [Ga21] and the references therein) and it is still a very active area of research [BiDeWe24, ChGaTh24, el24]. Most of the articles seem to be concerned with the properties that work exactly like in the $C^{*}$-case [el24] and a few others deal with new (perhaps surprising) phenomena [BiDeWe24, ChGaTh24]. In this sense, the present article should be put in the first group, as we plan to show that the topological stable rank is independent of $p$, for a large class of symmetrized crossed products. Symmetrized $L^{p}$-crossed products are ‘-analogs’ of regular $L^{p}$-crossed products and have the advantage of being Banach -algebras naturally acting on a natural $L^{2}$-space. They have been studied in [CP19, SaWi20, AuOr22, el24b, el24, BaKw25], often under the name of ‘algebras of pseudofunctions’.
In the case of topological stable rank 1, we can also say something about regular $L^{p}$-crossed products. In fact, we also aim to solve the following question of Christopher Phillips [CP13, Question 8.4].
Question 1.1.
Let $\alpha$ be a minimal action of $\mathbb{Z}$ either on the Cantor set $X=2^{\mathbb{N}}$ or on $X=\mathbb{T}$ by irrational rotations. Does $F^{p}(\mathbb{Z},X,\alpha)$ have topological stable rank 1?
Instead of $C^{*}$-dynamical systems, we chose to work with Fell bundles over groups of subexponential growth. This is in order to maximize generality, but also because the main tools we used were already developed in this framework [BeCo24, Fl24]. Note that the inclusion of Fell bundles provides almost no extra complication (when the approach is compared to, say, an approach based strictly on twisted $C^{*}$-dynamical systems). We also hope that this generality will be useful in the future.
[…]
Theorem 1.2.
[…]
${\rm tsr}(\mathfrak{B})={\rm tsr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}\quad\text{ and }\quad{\rm rr}(\mathfrak{B})={\rm rr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}.$
Corollary 1.3 (Corollary 2.11).
[…]
The class of Banach -algebras sitting between $\ell^{1}(G\,|\,{\mathscr{C}})$ and ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ seems to be vast. Indeed, for twisted $C^{*}$-dynamical systems of the form $(G,C_{0}(X),\alpha,\omega)$ and taking a subset $P\subset[1,\infty]$ as an input, Bardadyn and Kwaśniewski constructed an ‘$L^{P}$-completion’ of $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$. If $P$ is symmetric with respect to $2$, this is a Banach -algebra [BaKw25, Remark 4.13], which embeds in $C_{0}(X)\rtimes_{\alpha,\omega}G$ by the Riesz-Thorin interpolation theorem.
In particular, applying the above-mentioned theorem to symmetrized twisted $L^{p}$-crossed products yields the following theorem. In it, we use the notations $F^{p}_{*}(G,X,\alpha,\omega)$ and $F^{p}(G,X,\alpha,\omega)$ for the symmetrized and regular twisted $L^{p}$-crossed products associated with a twisted $C^{*}$-dynamical system of the form $(G,C_{0}(X),\alpha,\omega)$.
Theorem 1.4.
Let $(G,C_{0}(X),\alpha,\omega)$ be a twisted $C^{*}$-dynamical system, where $X$ is a Hausdorff locally compact space and $G$ is a group of subexponential growth. Then
${\rm rr}\big{(}F^{q}_{*}(G,X,\alpha,\omega)\big{)}={\rm rr}\big{(}C_{0}(X)\rtimes_{\alpha,\omega}G\big{)}\quad\text{ and }\quad{\rm tsr}\big{(}F^{q}_{*}(G,X,\alpha,\omega)\big{)}={\rm tsr}\big{(}C_{0}(X)\rtimes_{\alpha,\omega}G\big{)},$
for all $q\in[1,2]$. Furthermore, if ${\rm tsr}\big{(}C_{0}(X)\rtimes_{\alpha,\omega}G\big{)}=1$, then one also has
${\rm tsr}\big{(}F^{p}(G,X,\alpha,\omega)\big{)}=1,\quad\text{ for all }p\in[1,\infty].$
In particular, we have the following result. It fully settles Question 1.1.
Corollary 1.5.
Let $\alpha$ be a minimal action of $\mathbb{Z}$ either on the Cantor set $X=2^{\mathbb{N}}$ or on $X=\mathbb{T}$ by irrational rotations. Then $F^{p}(\mathbb{Z},X,\alpha)$ has topological stable rank $1$, for all $p\in[1,\infty]$.
In fact, and thanks to the generality of our setting, a lot of other topological ranks are computed in Theorem 3.2. They might be regarded as $L^{p}$-versions of results in [Pu90b, Pu90, BlKuRo92, Ka93, CP05]. It was recently announced by Li and Niu that every crossed product $C(X)\rtimes\mathbb{Z}^{d}$, coming from a free and minimal action $\mathbb{Z}^{d}\curvearrowright X$, has topological stable rank $1$ [LiNi23]. This widens the class of algebras where our theorems are applicable and provides other examples not explicitly mentioned here.
The reader should notice that some of the methods used in these papers are very different and some results are very intricate, but we succeeded in bypassing these methods.
The paper is organized as follows. Section 2 deals with preliminaries but also the proof of Theorem 1.2. It first introduces Fell bundles and their $\ell^{1}$-algebras and recalls some of the tools that will be later involved in the proofs of the main results. This section is then divided into two subsections. The first subsection introduces and studies the topological stable rank for these Fell bundle algebras, while the second subsection does the same for the real rank. Section 3 then briefly discusses the construction of $L^{p}$-crossed products and symmetrized $L^{p}$-crossed products and explains how the previously obtained results apply to these algebras. The main theorem of this section is Theorem 3.2, which basically consists of a list of results that follow from Theorem 3 and their (known) appropriate $C^{*}$-counterparts.
2 $\ell^{1}$-algebras of Fell bundles
From now on, $G$ will denote an amenable group with unit ${\sf e}$. We will also fix a Fell bundle $\mathscr{C}\!=\bigsqcup_{x\in G}\mathfrak{C}_{x}$ over $G$. The associated algebra of summable cross-sections $\ell^{1}(G\,|\,\mathscr{C})$ is a Banach -algebra and a completion of the space $C_{\rm c}(G\,|\,\mathscr{C})$ of cross-sections with finite support under the norm
$\lVert\Phi\rVert_{1}=\sum_{x\in G}\lVert\Phi(x)\rVert_{\mathfrak{C}_{x}}.$
The universal $C^{*}$-algebra completion of $\ell^{1}(G\,|\,\mathscr{C})$ is denoted by ${\rm C^{*}}(G\,|\,\mathscr{C})$ and the associated universal $C^{*}$-norm is denoted $\lVert\cdot\rVert_{*}$. For the general theory of Fell bundles, we recommend [FD88, Chapter VIII]. Here, we only recall the product on $\ell^{1}(G\,|\,\mathscr{C})$, given by
$\big{(}\Phi*\Psi\big{)}(x)=\sum_{y\in G}\Phi(y)\bullet\Psi(y^{-1}x)$
and its involution
$\Phi^{*}(x)=\Phi(x^{-1})^{\bullet}\,,$
in terms of the operations $\big{(}\bullet,^{\bullet}\big{)}$ on the Fell bundle. Note that the adjoint operation $\Phi\mapsto\Phi^{*}$ is isometric.
We will also consider the $\ell^{\infty}$-norm
$\lVert\Phi\rVert_{\infty}=\sup_{x\in G}\lVert\Phi(x)\rVert_{\mathfrak{C}_{x}},$
primarily because $\lVert\Phi\rVert_{\infty}\leq\lVert\Phi\rVert_{*}$, for all $\Phi\in{\rm C^{*}}(G\,|\,\mathscr{C})$ [Fl24, Lemma 3.4].
One of the main classes of examples of Fell bundles is presented next. It is the class of Fell bundles coming from twisted $C^{*}$-dynamical systems.
Example 2.1.
Let $\mathfrak{A}$ be a $C^{*}$-algebra. A (continuous) twisted action of $G$ on $\mathfrak{A}$ is a pair $(\alpha,\omega)$ of maps $\alpha:G\to{\rm Aut}({\mathfrak{A}})$, $\omega:G\times G\to\mathcal{UM}({\mathfrak{A}})$, such that
  • (i)
    $\alpha_{x}(\omega(y,z))\omega(x,yz)=\omega(x,y)\omega(xy,z)$,
  • (ii)
    $\alpha_{x}\big{(}\alpha_{y}(a)\big{)}\omega(x,y)=\omega(x,y)\alpha_{xy}(a)$,
  • (iii)
    $\omega(x,{\sf e})=\omega({\sf e},y)=1,\alpha_{\sf e}={\rm id}_{{\mathfrak{A}}}$,
for all $x,y,z\in G$ and $a\in\mathfrak{A}$.
The quadruple $(G,\mathfrak{A},\alpha,\omega)$ is called a twisted $C^{*}$-dynamical system. Given such a twisted action, one usually forms the so-called twisted convolution algebra $\ell^{1}_{\alpha,\omega}(G,\mathfrak{A})$, consisting of all summable sequences $\Phi:G\to\mathfrak{A}$, endowed with the product
$\Phi*\Psi(x)=\sum_{y\in G}\Phi(y)\alpha_{y}[\Psi(y^{-1}x)]\omega(y,y^{-1}x)$
and the involution
$\Phi^{*}(x)=\omega(x,x^{-1})^{*}\alpha_{x}[\Phi(x^{-1})^{*}],$
making $\ell^{1}_{\alpha,\omega}(G,\mathfrak{A})$ a Banach -algebra under the norm $\lVert\Phi\rVert_{1}=\sum_{x\in G}\lVert\Phi(x)\rVert_{\mathfrak{A}}$. When the twist is trivial ($\omega\equiv 1$), we omit any mention of it and call the resulting algebra $\ell^{1}_{\alpha}(G,\mathfrak{A})$ as the convolution algebra associated with the action $\alpha$. In this case, the triple $(G,\mathfrak{A},\alpha)$ is called a (untwisted) $C^{*}$-dynamical system. The associated $C^{*}$-envelope is usually denoted $\mathfrak{A}\rtimes_{\alpha,\omega}G$, or $\mathfrak{A}\rtimes_{\alpha}G$ in the untwisted case.
It is well known that these algebras correspond to our algebras $\ell^{1}(G\,|\,\mathscr{C})$, for very particular Fell bundles. In fact, these bundles may be easily described as $\mathscr{C}=\mathfrak{A}\times G$, with quotient map $q(a,x)=x$, constant norms $\lVert\cdot\rVert_{\mathfrak{C}_{x}}=\lVert\cdot\rVert_{\mathfrak{A}}$, and operations
$(a,x)\bullet(b,y)=(a\alpha_{x}(b)\omega(x,y),xy)\quad\textup{and}\quad(a,x)^{\bullet}=(\alpha_{x}^{-1}(a^{*})\omega(x^{-1},x)^{*},x^{-1}).$
In what follows, $\widetilde{\mathfrak{B}}$ will denote the minimal unitization of $\mathfrak{B}$ (if $\mathfrak{B}$ was already unital, we simply set $\widetilde{\mathfrak{B}}=\mathfrak{B}$). We note that the minimal unitization of $\ell^{1}(G\,|\,\mathscr{C})$ corresponds to the $\ell^{1}$-algebra of the ‘unitization Fell bundle’ $\widetilde{\mathscr{C}}\to G$, which is constructed from $\mathscr{C}$ simply by replacing $\mathfrak{C}_{\sf e}$ with $\widetilde{\mathfrak{C}_{\sf e}}$ and keeping all the other fibers intact.
In order to prove our main result, we first need to recall some notions of multipliers. In particular, the following is a well-known characterization of amenability that can be found in [BrOz08, Theorem 2.6.8]. It uses the notion of positive definite functions, so we recall that first.
Definition 2.2.
A function $\varphi:G\to\mathbb{C}$ is said to be positive definite if the matrix
$[\varphi(x^{-1}y)]_{x,y\in K}\in\mathbb{M}_{|K|}(\mathbb{C})$
is positive definite for every finite subset $K\subset G$.
Lemma 2.3.
$G$ is amenable if and only if there exists a net $\{\varphi_{j}\}_{j\in J}$ of finitely supported positive definite functions on $G$ such that $\varphi_{j}\to 1$ pointwise and $\varphi_{i}({\sf e})\leq 1$, for all $j\in J$.
Furthermore, Bédos and Conti [BeCo24, Example 4.5] showed that a positive definite function $\varphi:G\to\mathbb{C}$ induces a c.p. map
$T_{\varphi}:{\rm C^{*}}(G\,|\,\mathscr{C})\to{\rm C^{*}}(G\,|\,\mathscr{C}),\quad\text{ given by }\quad T_{\varphi}(\Phi)(x)=\varphi(x)\Phi(x).$
In fact, it is easy to see that, if $\varphi_{j}\to 1$ pointwise and $\varphi_{j}({\sf e})\leq 1$ (as in the previous lemma), then $\lVert T_{\varphi_{j}}\rVert\leq 1$ and $T_{\varphi_{j}}\Phi\to\Phi$, for all $\Phi\in C_{\rm c}(G\,|\,\mathscr{C})$.
The other ingredient we need to introduce comes from spectral theory. So, as usual, we will denote the spectrum of an element $b\in\mathfrak{B}$ by
${\rm Spec}_{\mathfrak{B}}(b)=\{\lambda\in\mathbb{C}\mid b-\lambda 1\textup{ isnot invertible in }\widetilde{\mathfrak{B}}\}.$
We also recall that a group is said to have subexponential growth if
$\lim_{n\to\infty}|K^{n}|^{1/n}=1,$
for all finite subsets $K\subset G$. All nilpotent groups have subexponential growth, so they serve as examples. In fact, all of our concrete applications are about nilpotent groups (see Section 3).
Theorem 2.4 ([Fl24]).
Suppose that $G$ has subexponential growth. Then
${\rm Spec}_{\ell^{1}(G\,|\,\mathscr{C})}(\Phi)={\rm Spec}_{{\rm C^{*}}(G\,|\,\mathscr{C})}(\Phi),\quad\text{ for all }\Phi\in C_{\rm c}(G\,|\,\mathscr{C}).$
Before moving to the next section, we remark that all the Banach -algebras mentioned in this section have a continuous (actually isometric) involution. In what follows, the algebras that admit a continuous involution will play a central role.
2.1  Topological stable rank
Let $\mathfrak{B}$ be a unital Banach algebra and $n\geq 1$. We set
${\rm Lg}_{n}(\mathfrak{B}):=\left\{(b_{1},\ldots,b_{n})\in\mathfrak{B}^{n}\mid\text{the left ideal generated by $b_{1},\ldots,b_{n}$ coincides with $\mathfrak{B}$}\right\}.$
The abbreviation ‘Lg’ stands for ‘left generators’. If $\mathfrak{B}$ has an involution, we also denote ${\rm Lg}_{n}(\mathfrak{B})_{\rm sa}:={\rm Lg}_{n}(\mathfrak{B})\cap\mathfrak{B}^{n}_{\rm sa},$ for the set of self-adjoint left generators. We now give the definition of (left) topological stable rank, as given by Rieffel in [Ri83, Definition 1.4].
Definition 2.5.
The (left) topological stable rank of a unital Banach algebra $\mathfrak{B}$ is the smallest $n\in\mathbb{Z}^{+}$ with the property that ${\rm Lg}_{n}(\mathfrak{B})$ is dense in $\mathfrak{B}^{n}$ under the product topology. In such a case, we write ${\rm tsr}(\mathfrak{B})=n$. On the other hand, we say that ${\rm tsr}(\mathfrak{B})=\infty$ if no such $n$ exists.
Rieffel observed that one gets the same notion if we decide to work with right ideals on a Banach -algebra, at least when the involution is continuous [Ri83, Proposition 1.6]. All the results we will prove here also follow for the right-sided topological stable rank either by virtue of this observation or by analogy. In any case, let us now introduce a characterization for the tuples in ${\rm Lg}_{n}(\mathfrak{B})$ that is easier to work with. It is well known to experts, but we include it here for completeness (and to highlight the fact that it uses $C^{*}$-techniques). In fact, it is sometimes taken as the definition (cf. [Th21, pgs. 16-17]).
Proposition 2.6.
Let $\mathfrak{B}$ be a unital $C^{*}$-algebra. Then $(b_{1},\ldots,b_{n})\in{\rm Lg}_{n}(\mathfrak{B})$ if and only if $\sum_{i=1}^{n}b_{i}^{*}b_{i}$ is invertible.
Proof.
If $\sum_{i=1}^{n}b_{i}^{*}b_{i}$ is invertible, then the left ideal generated by $b_{1},\ldots,b_{n}$ contains an invertible element, and so it coincides with $\mathfrak{B}$. On the other hand, and for the sake of contradiction, suppose that $b_{1},\ldots,b_{n}$ generates $\mathfrak{B}$ as a left ideal, but $\sum_{i=1}^{n}b_{i}^{*}b_{i}$ is not invertible.
Since $b_{1},\ldots,b_{n}$ generates $\mathfrak{B}$, we can find $a_{1},\ldots,a_{n}\in\mathfrak{B}$ such that
$1=a_{1}b_{1}+\ldots+a_{n}b_{n}.$
Now let $\pi:\mathfrak{B}\to\mathbb{B}(\mathcal{H})$ be a faithful -representation. Then
$0\in{\rm Spec}_{\mathfrak{B}}\Big{(}\sum_{i=1}^{n}b_{i}^{*}b_{i}\Big{)}={\rm Spec}_{\mathbb{B}(\mathcal{H})}\Big{(}\sum_{i=1}^{n}\pi(b_{i}^{*}b_{i})\Big{)},$
so $S=\sum_{i=1}^{n}\pi(b_{i}^{*}b_{i})$ is a positive non-invertible operator on $\mathcal{H}$. Since $S$ is not bounded below, there exists a unit vector $\xi\in\mathcal{H}$ such that
$\langle S\xi,\xi\rangle<\frac{1}{M^{2}n^{2}},\quad\text{ where }M=\max_{1\leq i\leq n}\lVert\pi(a_{i})\rVert_{\mathbb{B}(\mathcal{H})}.$
But then, for every $1\leq i\leq n$, we have
$\lVert\pi(b_{i})\xi\rVert_{\mathcal{H}}^{2}=\langle\pi(b_{i}^{*}b_{i})\xi,\xi\rangle<\frac{1}{M^{2}n^{2}}$ (2.1)
and
[…]
which is indeed a contradiction. ∎
Lemma 2.7.
Let $\mathfrak{A}\subset\mathfrak{B}$ be a dense, continuous unital inclusion of Banach algebras. Then ${\rm tsr}(\mathfrak{B})\leq{\rm tsr}(\mathfrak{A}).$
Proof.
Let ${\rm tsr}(\mathfrak{A})=n<\infty$, $\varepsilon>0$ and $b_{i}\in\mathfrak{B}$, for $1\leq i\leq n$. Then there exist elements $a_{i},a^{\prime}_{i},c_{i}\in\mathfrak{A}$, $1\leq i\leq n$, such that $\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$, $\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$ and
$1=\sum_{i=1}^{n}c_{i}a_{i}.$
But then the left ideal generated by $a_{1},\ldots,a_{n}$ in $\mathfrak{B}$ contains $1$, so $(a_{1},\ldots,a_{n})\in{\rm Lg}_{n}(\mathfrak{B})$. Since
$\lVert a_{i}-b_{i}\rVert_{\mathfrak{B}}\leq\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}+\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon,$
we see that ${\rm Lg}_{n}(\mathfrak{B})$ is dense in $\mathfrak{B}^{n}$, so ${\rm tsr}(\mathfrak{B})\leq n$. ∎
We now prove the main theorem of the section. The rough idea is to use the characterization of ${\rm Lg}_{n}(\mathfrak{B})$ in terms of invertibility and use Theorem 2.4 to ensure that the inverses stay inside of $\ell^{1}(G\,|\,{\mathscr{C}})$. Lemma 2.3 then provides us with a way of upgrading the convergence in ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ to convergence in $\ell^{1}(G\,|\,{\mathscr{C}})$.
Theorem 2.8.
Suppose that $G$ has subexponential growth. Let $\mathfrak{B}$ be a Banach -algebra such that $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}\subset{\rm C^{*}}(G\,|\,{\mathscr{C}})$. Then ${\rm tsr}(\mathfrak{B})={\rm tsr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}.$
Proof.
It is enough to prove the unital case, as passing to the unitization still gives a Fell bundle algebra that satisfies $\ell^{1}(G\,|\,\widetilde{\mathscr{C}})\subset\widetilde{\mathfrak{B}}\subset{\rm C^{*}}(G\,|\,\widetilde{\mathscr{C}})$. Applying Lemma 2.7, we have
${\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}\leq{\rm tsr}(\mathfrak{B})\leq{\rm tsr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}.$
So we only need to prove that ${\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}\geq{\rm tsr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}$.
Indeed, assume that ${\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=n<\infty$, let $\varepsilon>0$ and let $\Phi_{1},\ldots,\Phi_{n}\in C_{\rm c}(G\,|\,\mathscr{C})$ with $M=\max_{1\leq i\leq n}\lVert\Phi_{i}\rVert_{1}$ and use Lemma 2.3 to find a finitely supported positive definite function $\varphi:G\to\mathbb{C}$ such that
$\lVert\Phi_{i}-T_{\varphi}\Phi_{i}\rVert_{1}<\frac{\varepsilon}{2},\quad\text{for all }1\leq i\leq n.$ (2.2)
Now we use the definition of topological stable rank to find $\Psi_{1},\ldots,\Psi_{n}\in{\rm C^{*}}(G\,|\,{\mathscr{C}})$ that generate ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ as a left ideal and such that
$\lVert\Psi_{i}-\Phi_{i}\rVert_{*}<\frac{\varepsilon}{4|{\rm Supp}(\varphi)|},\quad\text{ for all }1\leq i\leq n.$ (2.3)
Because of Proposition 2.6, $\sum_{i=1}^{n}\Psi_{i}^{*}*\Psi_{i}$ is invertible in ${\rm C^{*}}(G\,|\,\mathscr{C})$. We now show that $T_{\varphi}\Psi_{i}$ approximates $\Phi_{i}$ in the $\ell^{1}$-norm.
$\lVert T_{\varphi}\Psi_{i}-\Phi_{i}\rVert_{1}$ $\leq\lVert T_{\varphi}\Psi_{i}-T_{\varphi}\Phi_{i}\rVert_{1}+\lVert\Phi_{i}-T_{\varphi}\Phi_{i}\rVert_{1}$
(2.2)${\leq}\lVert\Psi_{i}-\Phi_{i}\rVert_{\infty}|{\rm Supp}(\varphi)|+\frac{\varepsilon}{2}$
$\leq\lVert\Psi_{i}-\Phi_{i}\rVert_{*}|{\rm Supp}(\varphi)|+\frac{\varepsilon}{2}$
<(2.3)${<}\frac{3}{4}\varepsilon<\varepsilon.$
In particular, we have
$\|T_{\varphi}\Psi_{i}\|_{1}\leq\|\Phi_{i}\|_{1}+\frac{3}{4}\varepsilon\quad\text{ and }\quad\|\Psi_{i}\|_{*}\leq\|\Phi_{i}\|_{*}+\frac{\varepsilon}{4}.$ (2.4)
We also note that
$\Big{\|}\sum_{i=1}^{n}\Psi_{i}^{*}*\Psi_{i}-\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}\Big{\|}_{*}$ $\leq\sum_{i=1}^{n}\lVert\Psi_{i}^{*}*\Psi_{i}-(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}\rVert_{*}$
$\leq\sum_{i=1}^{n}\lVert\Psi_{i}-T_{\varphi}\Psi_{i}\rVert_{*}(\lVert\Psi_{i}\rVert_{*}+\lVert T_{\varphi}\Psi_{i}\rVert_{*})$
(2.4)${\leq}(2\max_{1\leq i\leq n}\lVert\Phi_{i}\rVert_{1}+\varepsilon)\sum_{i=1}^{n}\lVert\Psi_{i}-T_{\varphi}\Psi_{i}\rVert_{*},$
where the norms inside the sum can be estimated as
$\lVert\Psi_{i}-T_{\varphi}\Psi_{i}\rVert_{*}$ $\leq\lVert\Psi_{i}-\Phi_{i}\rVert_{*}+\lVert\Phi_{i}-T_{\varphi}\Phi_{i}\rVert_{*}+\lVert T_{\varphi}\Phi_{i}-T_{\varphi}\Psi_{i}\rVert_{*}$
$\leq 2\lVert\Psi_{i}-\Phi_{i}\rVert_{*}+\lVert\Phi_{i}-T_{\varphi}\Phi_{i}\rVert_{1}$
[…]
Hence, returning to the main line of computation, we see that
$\Big{\|}\sum_{i=1}^{n}\Psi_{i}^{*}*\Psi_{i}-\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}\Big{\|}_{*}\leq n\varepsilon(2M+\varepsilon).$
So, if $\varepsilon$ is small enough, $\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}$ is close to $\sum_{i=1}^{n}\Psi_{i}^{*}*\Psi_{i}$, so it has to be invertible. But $\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}\in C_{\rm c}(G\,|\,\mathscr{C})$ and, applying by Theorem 2.4, we see that it is invertible in $\ell^{1}(G\,|\,\mathscr{C})$. Reasoning as before, we conclude that $T_{\varphi}\Psi_{1},\ldots,T_{\varphi}\Psi_{n}$ generates $\ell^{1}(G\,|\,\mathscr{C})$ as a left ideal. Hence $(\Phi_{1},\ldots,\Phi_{n})$ belongs to the closure of ${\rm Lg}_{n}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}$ and this proves that ${\rm tsr}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}\leq n$. ∎
Corollary 2.9.
Suppose that $G$ has subexponential growth. Then ${\rm tsr}(\ell^{1}(G\,|\,{\mathscr{C}})\big{)}={\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}.$
Simply specializing to the twisted $C^{*}$-dynamical system case (i.e. the setting of Example 2.1) allows us to write the following corollary.
Corollary 2.10.
Suppose that $G$ has subexponential growth and let $(G,\mathfrak{A},\alpha,\omega)$ be a twisted $C^{*}$-dynamical system. Then ${\rm tsr}(\ell^{1}_{\alpha,\omega}(G,\mathfrak{A})\big{)}={\rm tsr}\big{(}\mathfrak{A}\rtimes_{\alpha,\omega}G\big{)}.$
Finally, in the case that $\mathfrak{B}$ is only a Banach algebra, we can only guarantee the equality in the case of topological stable rank 1.
Corollary 2.11.
Suppose that $G$ has subexponential growth and that ${\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=1$. Let $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}$ be any continuous, dense inclusion into a Banach algebra $\mathfrak{B}$. If $\ell^{1}(G\,|\,{\mathscr{C}})$ is unital, we also assume that the inclusion $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}$ is unital. Then ${\rm tsr}(\mathfrak{B})=1.$
Proof.
Combining Theorem 2.8 with Lemma 2.7 yields
${\rm tsr}(\mathfrak{B})\leq{\rm tsr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}={\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=1,$
which proves the claim. ∎
2.2  Real rank
We now proceed to our study of the real rank. The reader will notice that the similarity of definitions allows our line of argumentation to stay essentially the same, provided that we take care of self-adjointness and the corresponding shift in indices.
Definition 2.12.
The real rank of a unital Banach -algebra $\mathfrak{B}$ is the smallest $n\in\mathbb{N}$ with the property that ${\rm Lg}_{n+1}(\mathfrak{B})_{\rm sa}$ is dense in $\mathfrak{B}^{n+1}_{\rm sa}$ under the product topology. In such a case, we write ${\rm rr}(\mathfrak{B})=n$. On the other hand, we say that ${\rm rr}(\mathfrak{B})=\infty$ if no such $n$ exists.
As in the case of the topological stable rank, we can provide a more manageable characterization, at least for $C^{*}$-algebras.
Proposition 2.13.
Let $\mathfrak{B}$ be a unital $C^{*}$-algebra. Then $(b_{1},\ldots,b_{n})\in{\rm Lg}_{n}(\mathfrak{B})_{\rm sa}$ if and only if $\sum_{i=1}^{n}b_{i}^{2}$ is invertible and $b_{i}=b_{i}^{*}$ for all $1\leq i\leq n$.
Proof.
Immediate application of Proposition 2.6. ∎
Lemma 2.14.
Let $\mathfrak{A}\subset\mathfrak{B}$ be a dense, continuous unital inclusion of Banach -algebras. If $\mathfrak{A}$ has a continuous involution, then ${\rm rr}(\mathfrak{B})\leq{\rm rr}(\mathfrak{A}).$
Proof.
The proof is the same as the proof of Lemma 2.7, but we include it here. Let ${\rm rr}(\mathfrak{A})=n<\infty$, $\varepsilon>0$ and $b_{i}\in\mathfrak{B}_{\rm sa}$, for $0\leq i\leq n$. Then there exists elements $a_{i},a^{\prime}_{i}\in\mathfrak{A}_{\rm sa}$, $c_{i}\in\mathfrak{A}$, with $0\leq i\leq n$, such that $\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$, $\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$ and
$1=\sum_{i=0}^{n}c_{i}a_{i}.$
But then the left ideal generated by $a_{0},\ldots,a_{n}$ coincides with $\mathfrak{B}$, so $(a_{0},\ldots,a_{n})\in{\rm Lg}_{n+1}(\mathfrak{B})_{\rm sa}$. Since
$\lVert a_{i}-b_{i}\rVert_{\mathfrak{B}}\leq\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}+\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon,$
we get that ${\rm Lg}_{n+1}(\mathfrak{B})_{\rm sa}$ is dense in $\mathfrak{B}^{n+1}$. ∎
We now prove the real rank analog of Theorem 2.8. The proof follows the same lines.
Theorem 2.15.
Suppose that $G$ has subexponential growth. Let $\mathfrak{B}$ be a Banach -algebra with continuous involution and such that $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}\subset{\rm C^{*}}(G\,|\,{\mathscr{C}})$. Then ${\rm rr}(\mathfrak{B})={\rm rr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}.$
Proof.
As before, it is enough to prove the unital case and, because of Lemma 2.14, we only need to prove that ${\rm rr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}\geq{\rm rr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}$.
Suppose that ${\rm rr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=n<\infty$. We let $\varepsilon>0$ and $\Phi_{0},\ldots,\Phi_{n}\in C_{\rm c}(G\,|\,\mathscr{C})_{\rm sa}$ with $M=\max_{0\leq i\leq n}\lVert\Phi_{i}\rVert_{1}$ and use Lemma 2.3 to find a finitely supported positive definite function $\varphi:G\to\mathbb{C}$ such that
$\lVert\Phi_{i}-T_{\varphi}\Phi_{i}\rVert_{1}<\frac{\varepsilon}{2},\quad\text{for all }0\leq i\leq n.$
Now we use the definition of real rank to find $\Psi_{0},\ldots,\Psi_{n}\in{\rm C^{*}}(G\,|\,{\mathscr{C}})_{\rm sa}$ that generate ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ as a left ideal and such that
$\lVert\Psi_{i}-\Phi_{i}\rVert_{*}<\frac{\varepsilon}{4|{\rm Supp}(\varphi)|},\quad\text{ for all }0\leq i\leq n.$
By Lemma 2.13, $\sum_{i=0}^{n}\Psi_{i}^{2}$ is invertible in ${\rm C^{*}}(G\,|\,\mathscr{C})$. Furthermore, by the same arguments as in the proof of Theorem 2.8, we have
$\lVert T_{\varphi}\Psi_{i}-\Phi_{i}\rVert_{1}<\varepsilon\quad\text{ and }\quad\Big{\|}\sum_{i=0}^{n}\Psi_{i}^{2}-\sum_{i=0}^{n}(T_{\varphi}\Psi_{i})^{2}\Big{\|}_{*}\leq\varepsilon(n+1)(2M+\varepsilon).$
For small $\varepsilon$, $\sum_{i=0}^{n}(T_{\varphi}\Psi_{i})^{2}$ will be close to $\sum_{i=0}^{n}\Psi_{i}^{2}$, so it will have to be invertible. But $\sum_{i=0}^{n}(T_{\varphi}\Psi_{i})^{2}\in C_{\rm c}(G\,|\,\mathscr{C})$, so, again by Theorem 2.4, the inverse lies in $\ell^{1}(G\,|\,\mathscr{C})$. Because $\varphi$ is positive definite, this time $(T_{\varphi}\Psi_{0},\ldots,T_{\varphi}\Psi_{n})\in{\rm Lg}_{n+1}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}_{\rm sa}$ and thus ${\rm rr}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}\leq n$. ∎
Corollary 2.16.
Suppose that $G$ has subexponential growth. Then ${\rm rr}(\ell^{1}(G\,|\,{\mathscr{C}})\big{)}={\rm rr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}.$
To finish the section, we specialize the last corollary to the twisted $C^{*}$-dynamical system setting.
Corollary 2.17.
Suppose that $G$ has subexponential growth and let $(G,\mathfrak{A},\alpha,\omega)$ be a twisted $C^{*}$-dynamical system. Then ${\rm rr}(\ell^{1}_{\alpha,\omega}(G,\mathfrak{A})\big{)}={\rm rr}\big{(}\mathfrak{A}\rtimes_{\alpha,\omega}G\big{)}.$
3 Applications to symmetrized twisted $L^{p}$-crossed products
As mentioned before, the purpose of this section is to apply the previous results to the case of symmetrized $L^{p}$-crossed products. For that purpose, we will fix a twisted $C^{*}$-dynamical system $(G,C_{0}(X),\alpha,\omega)$, where $G$ is a group of subexponential growth, $X$ a Hausdorff locally compact space, $\alpha$ an action of $G$ on $X$ and $\omega:G\times G\to\mathcal{U}(C_{b}(X))$ a $2$-cocycle. Here, we abuse the notation and call $\alpha$ both the action on $X$ and the induced action on $C_{0}(X)$. That is
$\alpha_{x}(f)(z)=f\big{(}\alpha_{x^{-1}}(z)),$
where $x\in G,z\in X,f\in C_{0}(X)$.
It is well known that if $\mathfrak{B}$ has a contractive approximate unit and $\pi:\mathfrak{B}\to\mathbb{B}({\mathscr{E}})$ is a non-degenerate, contractive representation, then $\pi$ extends uniquely to a (unital) representation $\pi:\mathcal{M}(\mathfrak{B})\to\mathbb{B}({\mathscr{E}})$ of the multiplier algebra of $B$. This is always the case for $C_{0}(X)$.
Definition 3.1.
Let $p\in[1,\infty],$ and let $(G,C_{0}(X),\alpha,\omega)$ be a twisted $C^{*}$-dynamical system as above. Let $(Y,\mathcal{B},\mu)$ be a measure space. A covariant representation of $(G,C_{0}(X),\alpha,\omega)$ on $L^{p}(Y,\mu)$ is a pair $(v,\pi)$ consisting of a map $g\mapsto v_{g}$ from $G$ to the invertible operators on $L^{p}(Y,\mu)$ and a non-degenerate, contractive representation $\pi:C_{0}(X)\to\mathbb{B}(L^{p}(Y,\mu)),$ such that
$v_{x}v_{y}={\pi}(\omega(x,y))v_{xy},\qquad v_{x}\pi(f)v_{x}^{-1}=\pi(\alpha_{x}(f)),$
for all $x,y\in G$ and $f\in C_{0}(X)$.
Given a covariant representation $(v,\pi)$ on $L^{p}(Y,\mu)$, we define the associated integrated representation $\pi\rtimes v:\ell^{1}_{\alpha,\omega}(G,C_{0}(X))\to\mathbb{B}(L^{p}(Y,\mu))$ by
$(\pi\rtimes v)(\Phi)(\xi)=\sum_{x\in G}\pi(\Phi(x))(v_{x}(\xi)),$
for all $\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ and all $\xi\in L^{p}(Y,\mu)$.
It is well known that every representation of $C_{0}(X)$ induces a covariant representation of $(G,C_{0}(X),\alpha,\omega)$. We proceed to describe this process. Indeed, let $\pi_{0}:C_{0}(X)\to\mathbb{B}(L^{p}(Y,\mu))$ be a contractive representation and let $\nu$ denote the counting measure on $G$. Then the following formulas define representations $v$ of $G$ and $\pi$ of $C_{0}(X)$ on $\ell^{p}(G,\,L^{p}(Y,\mu))\cong L^{p}(G\times Y,\nu\times\mu)$:
$v_{y}(\xi)(x)=\pi_{0}(\alpha_{x}^{-1}(\omega(y,y^{-1}x)))\xi(y^{-1}x),\quad\text{ and }\quad\big{(}\pi(f)\xi\big{)}(x)=\pi_{0}(\alpha_{x}^{-1}(f))\big{(}\xi(x)\big{)},$
for all $x,y\in G$, $\xi\in\ell^{p}(G,\,L^{p}(Y,\mu))$, and $f\in C_{0}(X)$.
It is easy to see that the pair $(v,\pi)$ constructed above is a covariant representation, where $v$ is isometric and $\pi$ contractive. It is usually called the regular covariant representation associated with $\pi_{0}$ [CP13, Definition 2.12]. If $\pi_{0}$ is the point-evaluation representation
$\pi_{0}:C_{0}(X)\to\mathbb{C},\quad\text{ given by }\pi_{0}(f)=f(z),$
then we use the notation $(v,\pi_{z})$ for the associated regular covariant representation. In that case, the integrated representation acquires the following form:
$(\pi_{z}\rtimes v)(\Phi)\xi(x)=\sum_{y\in G}\Phi(y)(xz)\xi(y^{-1}x)\omega(y,y^{-1}x)(xz),$
for all $x\in G,\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ and all $\xi\in\ell^{p}(G)$.
With these definitions at hand, we can introduce the following norms on $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$.
$\|\Phi\|_{p,\lambda}$ $=\sup\{\|(\pi\rtimes v)(\Phi)\|\mid(v,\pi)\text{ is a regular covariant pair}\},$
$\|\Phi\|_{p,r}$ $=\sup\{\|(\pi_{z}\rtimes v)(\Phi)\|\mid(v,\pi_{z})\text{ is a regular covariant pair and }z\in X\}.$
The completion $F^{p}(G,X,\alpha,\omega):=\overline{\ell^{1}_{\alpha,\omega}(G,C_{0}(X))}^{\|\cdot\|_{p,\lambda}}$ is usually called the reduced $L^{p}$-crossed product associated with $(G,C_{0}(X),\alpha,\omega)$. It is easily seen to be a Banach algebra. There is also a ‘full’ crossed product construction, which is obtained using all representations into $L^{p}$-spaces and not just the regular ones, but in our setting (the groups are amenable), these two constructions are known to coincide [BaKw25, Theorem 3.5], so we avoid introducing it.
On the other hand, the norms $\|\cdot\|_{p,r}$ can and have been used to introduce symmetrized versions of the $L^{p}$-crossed product (cf. [AuOr22]). Indeed, we let $p,q\in[1,\infty]$ be Hölder duals and define $F^{q}_{*}(G,X,\alpha,\omega)$ to be the completion of $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ under the norm
$\lVert\Phi\rVert_{q,*}=\max\{\|\Phi\|_{p,r},\|\Phi\|_{q,r}\}.$
The point is that $F^{q}_{*}(G,X,\alpha,\omega)$ is naturally a Banach -algebra that sits between $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ and $C_{0}(X)\rtimes_{\alpha,\omega}G$, so it fits our setting perfectly. Indeed, the fact that this algebra has a continuous involution comes from the fact that, for all $\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X)),\xi\in\ell^{p}(G),\eta\in\ell^{q}(G)$, one has
$\langle(\pi_{z}\rtimes v)(\Phi)\xi,\eta\rangle=\langle\xi,(\pi_{z}\rtimes v)(\Phi^{*})\eta\rangle,$
[…]
We remit the reader to [AuOr22, Section 3] or [BaKw25, Sections 2,3] for a careful discussion of the algebra $F^{q}_{*}(G,X,\alpha,\omega)$ and of these properties that we just mentioned. In any case, Theorem 1.2 applies to $F^{q}_{*}(G,X,\alpha,\omega)$ and we immediately obtain Theorem 1.4 as a consequence. The next theorem compiles consequences and explicit computations derived from these theorems.
Theorem 3.2.
The following are true.
  • (i)
    Let $G$ be an abelian group, whose Pontryagin dual is denoted $\widehat{G}$. Then, for $p\in[1,2]$,
    ${\rm rr}\big{(}F^{p}_{*}(G)\big{)}={\rm dim}\big{(}\widehat{G}\big{)}\quad\text{ and }\quad{\rm tsr}\big{(}F^{p}_{*}(G)\big{)}=\lfloor\tfrac{1}{2}{\rm dim}\big{(}\widehat{G}\big{)}\rfloor+1,$
    where ${\rm dim}(X)$ is the covering dimension of a compact space $X$.
  • (ii)
    Let $G$ be a nilpotent group and let $G^{f}\subset G$ denote the (normal) subgroup of torsion elements. Then, for $p\in[1,2]$,
    1. 1.
      ${\rm rr}\big{(}F^{p}_{*}(G)\big{)}=0$ if and only if $G=G^{f}$.
    2. 2.
      ${\rm rr}\big{(}F^{p}_{*}(G)\big{)}=1$ if and only if $G/G^{f}\cong\mathbb{Z}$.
  • (iii)
    Let $\alpha$ be a free, minimal action of $\mathbb{Z}^{d}$ on the Cantor set $X=2^{\mathbb{N}}$. Then
    ${\rm rr}\big{(}F^{q}_{*}(\mathbb{Z}^{d},X,\alpha)\big{)}=0\quad\text{ and }\quad{\rm tsr}\big{(}F^{q}_{*}(\mathbb{Z}^{d},X,\alpha)\big{)}={\rm tsr}\big{(}F^{p}(\mathbb{Z}^{d},X,\alpha)\big{)}=1,$
    for all $p\in[1,\infty)$ and $q\in[1,2]$.
  • (iv)
    Let $\alpha$ be an action of $\mathbb{Z}$ on $\mathbb{T}$ by irrational rotations. Then
    ${\rm rr}\big{(}F^{q}_{*}(\mathbb{Z},\mathbb{T},\alpha)\big{)}=0\quad\text{ and}\quad{\rm tsr}\big{(}F^{q}_{*}(\mathbb{Z},\mathbb{T},\alpha)\big{)}={\rm tsr}\big{(}F^{p}(\mathbb{Z},\mathbb{T},\alpha)\big{)}=1,$
    for all $p\in[1,\infty)$ and $q\in[1,2]$.
  • (v)
    Let $G$ be a group of subexponential growth and let $\alpha$ be the canonical action of $G$ on the Stone-Čech compactification $\beta G$. Then $G$ is locally finite if and only if ${\rm tsr}(F^{q}_{*}(G,\beta G,\alpha))=1$ for some (and hence any) $q\in[1,2]$.
Proof.
All the affirmations follow from Theorem 1.2, Corollary 2.11 and the fact that the appropriate $C^{*}$-counterpart is true. Indeed, (i) follows from [Ri83, BP91], (ii) follows from [Ka93, Theorem 2], (iii) follows from [CP05], (iv) follows from [Pu90, BlKuRo92] and (v) follows from [LiLi18, Corollary 1.3]. ∎
Acknowledgments
The author thankfully acknowledges support by the NSF grant DMS-2144739. He is grateful to Professors Ben Hayes and Hannes Thiel for their helpful comments and to Iason Moutzouris, who pointed him to Proposition 2.6. The author also thanks Professors Jianguo Zhang and Bartosz Kwaśniewski, who pointed him to the references [LiNi23] and [BaKw25], respectively.
[…]
ADDRESS
Felipe I. Flores
Department of Mathematics, University of Virginia,
114 Kerchof Hall. 141 Cabell Dr,
Charlottesville, Virginia, United States
E-mail: hmy3tf@virginia.edu